A student takes out a five-year loan of 1000. interest on the loan is at an annual effective interest rate of i. at the end of each year, the student pays the interest due on the loan and makes a deposit of twice the amount of that interest payment into a sinking fund. the sinking fund credits interest at an annual effective rate of 0.8i. the sinking fund will accumulate the amount needed to pay off the loan at the end of five years. calculate i.
Answer #1

Amount of interest payment each year=1000*i

Amount of sinking fund deposit=1000*i*2

Future value of sinking fund deposit=Loan amount

1000*i*2/(0.8*i)*((1+0.8*i)^5-1)=1000
=>(1+0.8*i)^5-1=0.4
=>i=((1.4^(1/5))-1)/0.8
=>i=8.7013%